OpenStudy (anonymous):

Evaluate: a^2 -3 (b^2-a^3)-b^0 if a= -3 and b= -5 a. -148 b. 14 c. -16 d. -142 e. none of the above

OpenStudy (anonymous):

i got none of these as my answer

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Plug in -3 anytime you see a and -5 anytime you see b

OpenStudy (anonymous):

(-3)^2-3(-5^2--3^3)-(-5)^0

OpenStudy (anonymous):

(9)-3(2)-1 then distribute 9-6-1 =2 so e

OpenStudy (anonymous):

that's what i did. i got the same answer. so E was right then?

jimthompson5910 (jim_thompson5910):

nope it's not E

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

i thinks it e

OpenStudy (anonymous):

Oh, i got the same answer as @xlilly did. what would the right way to do this be?

jimthompson5910 (jim_thompson5910):

a^2 -3 (b^2-a^3)-b^0 (-3)^2 -3 (b^2-(-3)^3)-b^0 .. plug in a = -3 (-3)^2 -3 ((-5)^2-(-3)^3)-(-5)^0 .. plug in b = -5 9 -3 (25-(-27))-1 9 -3 (25+27)-1 9 -3 (52)-1 9 -156-1 -147-1 -148 So, a^2 -3 (b^2-a^3)-b^0 = -148 when a= -3 and b= -5